# A halo model approach for mock catalogs of time-variable strong gravitational lenses

### v4

@@ 2.5 Lensing calculations
The calculation of gravitational lensing effects is performed using the public code glafic (Oguri, 2010, 2021). For each host halo and subhalo, we solve the lens equation to derive multiple images, as well as magnifications and time delays for individual images. Here, for numerical reasons, we do not calculate the lens equation with all deflectors, including a host halo and the subhalos, simultaneously, but break them down into individual deflector systems. To summarize, the lensing calculation of each host halo involves three deflector components: the elliptical NFW profile for the host halo, the elliptical Hernquist profile for the central galaxy, and external shear, while the lensing calculation of each subhalo involves five deflector components: the elliptical NFW profile for the host halo, the elliptical Hernquist profile for the central galaxy, the elliptical NFW profile for the subhalo, the elliptical Hernquist profile for the satellite galaxy, and external shear. While we expect some correlation of external shear between a halo and a subhalo within the halo or between subhalos within a halo, we ignore such correlation for simplicity. This approximation not to calculate the lens equation with all deflectors would be validated for two reasons. Firstly, as described in Sec. 2, we adopt $M_{\mathrm{tot}}-M^{i}_{\mathrm{sub}}$ as the mass of the host halo when performing lensing calculations for $i$-th subhalos of mass $M^{i}_{\mathrm{sub}}$. Therefore, influences from other subhalos are collectively included as the influence from the host halo. Secondly, it would be very rare to have several subhalos located within a scale of about their Einstein radii.
@@
Symbol $X$ Explanation Mean $\bar{X}$ Variance $\sigma_{X\mathrm{or}\ln(X)}$ Distribution $M_{\mathrm{hh}}$ mass of host halo in units of $M_{\odot}$ Tinker et al. (2008) $M_{\mathrm{tot}}$ total mass of halo in units of $M_{\odot}$ Tinker et al. (2008) $M_{\mathrm{sh}}$ mass of subhalo in units of $M_{\odot}$  Bond et al. (1991); Bower (1991); Lacey and Cole (1993) $M_{\mathrm{h}}$ mass of subhalo at the accretion time in units of $M_{\odot}$ Oguri and Takahashi (2020) $c_{\mathrm{h}}$ concentration parameter for host halos DKJ 0.33 Log-normal dist. (Eq. (4)) $c_{\mathrm{sh}}$ concentration parameter for subhalos Eq. (19) 0.33 Log-normal dist. (Eq. (4)) $e_{\mathrm{h}}$ ellipticity for host halos Eq. (8) 0.13 Truncated normal dist. (see below Eq. (8)) $e_{\mathrm{sh}}$ ellipticity for subhalos Eq. (8) 0.13 Truncated normal dist. (see below Eq. (8)) $\phi_{\mathrm{h}}$ position angle for host halos Uniform in $[-180,180]$ deg $\phi_{\mathrm{sh}}$ position angle for subhalos Uniform in $[-180,180]$ deg $M_{\mathrm{c*}}$ stellar mass of central galaxy in units of $M_{\odot}$ Behroozi et al. (2019) 0.20 Log-normal dist. $M_{\mathrm{s*}}$ stellar mass of satellite galaxy in units of $M_{\odot}$ Behroozi et al. (2019) 0.20 Log-normal dist. $r_{e}$ effective radius of central/satellite galaxy in units of $h^{-1}\mathrm{kpc}$ Eq. (10) Eq. (11) Log-normal dist. $e_{\mathrm{c*}}$ ellipticity for central galaxies 0.30 0.16 Truncated normal dist. (see below Eq. (11)) $e_{\mathrm{s*}}$ ellipticity for satellite galaxies 0.30 0.16 Truncated normal dist. (see below Eq. (11)) $\phi_{\mathrm{c*}}$ position angle for central galaxies $\phi_{\mathrm{h}}$ 35.4 Normal dist. $\phi_{\mathrm{s*}}$ position angle for satellite galaxies $\phi_{\mathrm{sh}}$ 35.4 Normal dist. $\gamma_{\mathrm{ext,1}}$ external shear component 0 Eq. (22) Truncated normal dist. (Eq. (21)) $\gamma_{\mathrm{ext,2}}$ external shear component 0 Eq. (22) Truncated normal dist. (Eq. (21))
@@ 3.1 Velocity dispersion function
We calculate velocity dispersions of galaxies in our compound lens model using the GalKin package in the Lenstronomy python library (Birrer and Amara, 2018; Birrer et al., 2021). In order to derive the velocity dispersion function of galaxies in our model, we first specify the field of view to populate halos and galaxies and then create a mock catalog of halos and galaxies in that region for each redshift bin as described in Sec. 2. We include not only main halos and central galaxies but also subhalos and satellite galaxies. In the latter case, we adopt the masses at their accretion time for computing velocity dispersions because we use those masses for calculating lensing properties of subhalos (see the discussion in Sec. 2.3 for more details). We then use the GalKin package to compute the velocity dispersion for each galaxy in the mock catalog, assuming the NFW and Hernquist profiles as density profiles of dark matter and stellar mass components, respectively. Here, the effect of the ellipticity is ignored because, as explained later, velocity dispersions are obtained from observations after circularization.
We compare the velocity dispersion function predicted by our model with that from OM10+, which is based on Bernardi et al. (2010). In Bernardi et al. (2010), the velocity dispersion function derived from the SDSS Data Release 6 is fitted to the Schechter function (Sheth et al., 2003) as
@@
with $\varphi_{*}=2.099\times 10^{-2}(h/0.7)^{3}\mathrm{Mpc}^{-3},\sigma_{*}=113.78\mathrm{~{}km}\mathrm{~{}s}^{-1}$, $\alpha=0.94$, and $\beta=1.85$. OM10+ also takes account of the redshift evolution of the velocity dispersion function, mainly reflecting an evolution in the mass-to-light ratio of galaxies (Treu et al., 2005; van Dokkum and van der Marel, 2007; Saglia et al., 2010; Holden et al., 2010; van de Sande et al., 2014; Zahid et al., 2015), as
@@
Here $\varphi_{\text{hyd }}(\sigma,z)$ is the velocity dispersion function presented in Torrey et al. (2015), which is derived from the Illustris simulation (Vogelsberger et al., 2014a, b; Genel et al., 2014). We caution that Eqs. (23) and (24) are used only for the validation test conducted in this Section, and not used for our main calculations to generate mock catalogs based on the halo model approach.
It should be noted that this velocity dispersion function is for $\sigma_{e/8}$, i.e., a luminosity-weighted velocity dispersion averaged within the effective radius $r_{e}/8$, rather than within $r_{e}$ [^1] Conventionally, this luminosity-weighted velocity dispersion within $r=r_{e}/8$ has often been adopted, historically based on the aperture size of previous observations. In this study, since we are considering strong gravitational lensing effects and the Einstein radii are typically comparable to $r=r_{e}$ (e.g., Oguri et al., 2014), the luminosity-weighted velocity dispersion averaged within $r=r_{e}$ is more relevant than $r=r_{e}/8$.. Hence, we convert the velocity dispersion function of $\sigma_{e/8}$ to that of $\sigma_{e}$ using the commonly-used relation for elliptical galaxies, which is derived from 40 early-type galaxies of the SAURON integral field spectroscopy data in Cappellari et al. (2006),
@@
Fig. 2 shows the comparison between the velocity dispersion function derived from our model and that of OM10+ calculated from Eq. (24) for several redshifts. For numerical reasons, we set a lower limit on the mass of halos to be calculated, $M_{\mathrm{hh}}\gtrsim 10^{11}~{}M_{\odot}$ (or $M_{\mathrm{sh}}\gtrsim 10^{10}~{}M_{\odot}$ for subhalos), which may lead to the underestimation of the number of galaxies with low velocity dispersions in our model. We find that both the velocity dispersion functions agree well for low redshifts ($z\lesssim 0.5$), which are main redshifts of lensing objects for lensed QSOs and SNe. In contrast, we find some mismatch at high redshifts ($z\gtrsim 0.5$), indicating that the redshift evolution in our model deviates from that in the Illustris simulation. To be more quantitative, we conduct the Kolmogorov-Smirnov test focusing on the velocity dispersion function at high velocity dispersions $v\geq 200\mathrm{~{}km}\mathrm{~{}s}^{-1}$ to find that p-values are larger than 0.002 for $z\leq 0.25$ and are smaller than $2\times 10^{-5}$ for $z\geq 0.5$, indicating that the mismatch of two velocity dispersion functions is indeed significant at high redshifts. Although the Illustris simulations were calibrated to match a set of global galaxy properties, they predicted galaxy sizes about twice as large as those observed (Genel et al., 2014). In the updated IllustrisTNG simulations incorporating AGN feedback, the predicted galaxy sizes are significantly reduced (Pillepich et al., 2018; Genel et al., 2018). Since there is a negative correlation between the galaxy size and the velocity dispersion for given stellar and halo masses, it is possible that the original Illustris simulations predict significantly smaller velocity dispersion at such high redshifts. As described in Sec. 2, our model adopts galaxy sizes directly measured from observations. Therefore, we expect our model to predict the redshift evolution of the velocity dispersion function more accurately and reliably than OM10+.
@@
Figure 2: Comparison between the velocity dispersion function derived with our model described in Sec. 2 (red) and that presented in Oguri (2018) (grey). The vertical axis shows the number of galaxies in each redshift bin with the width of $\Delta z=0.01$ for each velocity dispersion bin. Here, we set the area written on each panel to derive the number. Top left: $z=0.05$, Top center: $z=0.1$, Top right: $z=0.25$, Bottom left: $z=0.5$, Bottom center: $z=1.0$, Bottom right: $z=2.0$.
@@ 3.2 Effective radii and velocity dispersions at intermediate redshift
As mentioned above, our model of the effective radius of galaxies is based on measurements with JWST for galaxies at redshift $2.50.5<z<$ 2.5. For the intermediate redshift range of $0.2\lesssim z\lesssim 0.6$, there is a galaxy catalog from the SHELS survey of the F2 field of the Deep Lens Survey (Wittman et al., 2006; Geller et al., 2014; Damjanov et al., 2022). As an additional validity check, we compare effective radii and velocity dispersions for early-type galaxies at $0.2\lesssim z\lesssim 0.6$ in that galaxy catalog with those predicted by our model. The four panels on the left side of Fig. 3 show the effective radii of galaxies as a function of stellar mass, focusing on massive galaxies with masses of $M_{*}\gtrsim 10^{10.7}~{}M_{\odot}$ at $0.30.2<z<$ 0.3, $0.40.3<z<$ 0.4, $0.50.4<z<$ 0.5, and $0.60.5<z<$ 0.6. The right panel displays the corresponding velocity dispersions. For both effective radii and velocity dispersions, the SHELS F2 catalog galaxies, especially more massive galaxies than $10^{11}~{}M_{\odot}$, appear to have somewhat larger value than our mock galaxies, which is, however in agreement at the $2\sigma$ level.
@@
Figure 3: Comparison between our model with observation data of SHELS F2 at intermediate redshifts, $z\lesssim 0.6$. Only the galaxy samples with stellar masses of $M_{*}\gtrsim 10^{10.7}~{}M_{\odot}$ are plotted. The four panels on the left side show effective radii as a function of stellar mass at $0.30.2<z<$ 0.3, $0.40.3<z<$ 0.4, $0.50.4<z<$ 0.5, and $0.60.5<z<$ 0.6, while the four panels on the right represent those of velocity dispersions. The thick solid lines show the center lines for the galaxy samples. The dark red and light red regions show the $68\%$ and $95\%$ areas for our galaxy samples, while the thin grey solid and dotted lines show the $68\%$ and $95\%$ areas for SHELS F2 galaxy samples. We set the area of $4.2~{}\mathrm{deg}^{2}$ to create a mock galaxy sample, matching the area covered by the SHELS F2 catalog. Note that we do not plot the error bars for the effective radii and velocity dispersions for the SHELS F2 data.
@@ 3.3 Intrinsic scatter of velocity dispersions
Next, we discuss the intrinsic scatter of velocity dispersions, $\sigma_{e}$, for a given stellar mass of galaxies, which is also important in order to create a realistic mock catalog. For instance, ignoring the intrinsic scatter typically is suggested to lead to an underestimation of the slope of the $M_{*}$-$\sigma_{e}$ relation (see, e.g., Auger et al., 2010). Using the mock catalog obtained for the comparison of the velocity dispersion function, we can estimate the scatter of $\sigma_{e}$ at given $M_{*}$ and $z$. In the case of our model, the intrinsic scatter originates from three sources, (1) the scatter of the stellar-mass-halo-mass relation (Golden-Marx et al., 2022), (2) the scatter of the halo mass-concentration parameter relation (Macciò et al., 2007), and (3) the scatter of the galaxy size-mass relation (vdW24). We note that vdW24 claims that measurement uncertainties in observations are smaller than the intrinsic scatter of $M_{*}-r_{e}$ relation. Therefore, we assume that the scatter obtained from our model is also dominated by the intrinsic scatter.
@@
Cannarozzo et al. (2020) suggest that the intrinsic scatter of the $M_{*}$-$\sigma_{e}$ relation is about $0.075_{-0.003}^{+0.003}~{}\mathrm{dex}$ at given $M_{*}$ for early-type galaxies with masses $M_{*}\gtrsim 3.0\times 10^{10}~{}M_{\odot}$ in the SDSS sample. They also investigate the redshift evolution of the intrinsic scatter using the sample of the massive early-type galaxies at higher redshifts, $z\lesssim 2$, composed of 26 galaxies drawn from the LRIS sample presented in Belli et al. (2014), 56 galaxies from van de Sande et al. (2013), four galaxies from Gargiulo et al. (2015), and 24 galaxies from Belli et al. (2017). As a result, it is found that the redshift evolution of the scatter is weak in terms of the Bayesian factor comparison.
@@
We compute the intrinsic scatter in a stellar mass range of $11.710.5<\log_{10}(M_{*}/M_{\odot})<$ 11.7 at $z<2$. First, we divide our mock galaxies within this mass range into ($z,M_{*}$)-bins. We then calculate the scatter of velocity dispersion for the galaxy samples in each bin. The histograms of the scatter are plotted in Fig. 4, divided into three redshift ranges: $0.20.05<z<$ 0.2, $0.60.2<z<$ 0.6, and $2.00.6<z<$ 2.0. For the galaxy samples, we set $(\Delta z,\Delta\log_{10}(M*/M_{\odot}))=(0.01,0.02)$ for $0.20.05<z<$ 0.2, $(0.02,0.02)$ for $0.60.2<z<$ 0.6, and $(0.1,0.02)$ for $2.00.6<z<$ 2.0 for numerical reasons. It should be noted that the histograms in Fig. 4 include only the scatter from bins containing a sufficient number of galaxy samples, $N_{\mathrm{gal}}^{(z,M_{*})-\mathrm{bin}}>\mathcal{O}(10^{2})$. In the left panel of Fig. 4, we also plot the result of Cannarozzo et al. (2020) for SDSS samples containing galaxies at $0.20.05<z<$ 0.2 for comparison. We find that our model predicts the intrinsic scatter of about $0.072^{+0.002}_{-0.002}$ dex at $0.20.05<z<$ 0.2, which agrees with the result of Cannarozzo et al. (2020) at $\sim 1\sigma$ level. We also find that our model does not exhibit a strong redshift dependence on the intrinsic scatter. Since Cannarozzo et al. (2020) suggest the weak redshift evolution of the scatter, both results appear to be consistent.
@@
Figure 4: The scatter of $\sigma_{e}$ for mock galaxies at given $z$ and $M_{*}$ calculated with our model. We compute the scatter only for mock galaxies with $11.710.5<\log_{10}(M_{*}/M_{\odot})<$ 11.7 at $z<2$. We set $2000~{}\mathrm{deg}^{2}$, $100~{}\mathrm{deg}^{2}$, and $20~{}\mathrm{deg}^{2}$ to create mock galaxy samples at $0.20.05<z<$ 0.2 (left), $0.60.2<z<$ 0.6 (middle), and $2.00.6<z<$ 2.0 (right), respectively. The calculation method of the scatter is as follows. We first divide our mock galaxies into ($z,M_{*}$)-bins. We then calculate the scatter of velocity dispersion for the sample of galaxies in each bin. We finally plot these histograms only using those bins that hold more than $\mathcal{O}(10^{2})$ galaxy samples to ensure that the variance is properly computed. Here we set $(\Delta z,\Delta\log_{10}(M*/M_{\odot}))=(0.01,0.02)$ for the galaxy sample at $0.20.05<z<$ 0.2, $(0.02,0.02)$ at $0.60.2<z<$ 0.6, and $(0.1,0.02)$ at $2.00.6<z<$ 2.0 for a numerical reason. The grey dotted line appearing only in the left panel represents a best-fit value obtained for the SDSS galaxy sample at $0.20.05<z<$ 0.2 in Cannarozzo et al. (2020), $\sigma_{\ln\sigma_{e}}=0.075^{+0.003}_{-0.003}~{}\mathrm{dex}$. The orange-colored regions represent the ±$\pm 1\sigma$, ±$\pm 2\sigma$, and ±$\pm 3\sigma$ regions, in order from darkest to lightest.

### v5

@@ 2.5 Lensing calculations
The calculation of gravitational lensing effects is performed using the public code glafic (2010PASJ...62.1017O; 2021PASP..133g4504O). For each host halo and subhalo, we solve the lens equation to derive multiple images, as well as magnifications and time delays for individual images. Here, for numerical reasons, we do not calculate the lens equation with all deflectors, including a host halo and the subhalos, simultaneously, but break them down into individual deflector systems. To summarize, the lensing calculation of each host halo involves three deflector components: the elliptical NFW profile for the host halo, the elliptical Hernquist profile for the central galaxy, and external shear, while the lensing calculation of each subhalo involves five deflector components: the elliptical NFW profile for the host halo, the elliptical Hernquist profile for the central galaxy, the elliptical NFW profile for the subhalo, the elliptical Hernquist profile for the satellite galaxy, and external shear. While we expect some correlation of external shear between a halo and a subhalo within the halo or between subhalos within a halo, we ignore such correlation for simplicity. This approximation not to calculate the lens equation with all deflectors would be validated for two reasons. Firstly, as described in Sec. 2, we adopt $M_{\mathrm{tot}}-M^{i}_{\mathrm{sub}}$ as the mass of the host halo when performing lensing calculations for $i$-th subhalos of mass $M^{i}_{\mathrm{sub}}$. Therefore, influences from other subhalos are collectively included as the influence from the host halo. Secondly, it would be very rare to have several subhalos located within a scale of about their Einstein radii.
@@
Symbol $X$ Explanation Mean $\bar{X}$ Variance $\sigma_{X\mathrm{or}\ln(X)}$ Distribution $M_{\mathrm{hh}}$ mass of host halo in units of $M_{\odot}$ 2008ApJ...688..709T $M_{\mathrm{tot}}$ total mass of halo in units of $M_{\odot}$ 2008ApJ...688..709T $M_{\mathrm{sh}}$ mass of subhalo in units of $M_{\odot}$  1991ApJ...379..440B; 1991MNRAS.248..332B; 1993MNRAS.262..627L $M_{\mathrm{h}}$ mass of subhalo at the accretion time in units of $M_{\odot}$ 2020ApJ...901...58O $c_{\mathrm{h}}$ concentration parameter for host halos DKJ 0.33 Log-normal dist. (Eq. (4)) $c_{\mathrm{sh}}$ concentration parameter for subhalos Eq. (19) 0.33 Log-normal dist. (Eq. (4)) $e_{\mathrm{h}}$ ellipticity for host halos Eq. (8) 0.13 Truncated normal dist. (see below Eq. (8)) $e_{\mathrm{sh}}$ ellipticity for subhalos Eq. (8) 0.13 Truncated normal dist. (see below Eq. (8)) $\phi_{\mathrm{h}}$ position angle for host halos Uniform in [180,180][-180,180] deg $\phi_{\mathrm{sh}}$ position angle for subhalos Uniform in [180,180][-180,180] deg $M_{\mathrm{c*}}$ stellar mass of central galaxy in units of $M_{\odot}$ 2019MNRAS.488.3143B 0.20 Log-normal dist. $M_{\mathrm{s*}}$ stellar mass of satellite galaxy in units of $M_{\odot}$ 2019MNRAS.488.3143B 0.20 Log-normal dist. $r_{e}$ effective radius of central/satellite galaxy in units of $h^{-1}\mathrm{kpc}$ Eq. (10) Eq. (11) Log-normal dist. $e_{\mathrm{c*}}$ ellipticity for central galaxies 0.30 0.16 Truncated normal dist. (see below Eq. (11)) $e_{\mathrm{s*}}$ ellipticity for satellite galaxies 0.30 0.16 Truncated normal dist. (see below Eq. (11)) $\phi_{\mathrm{c*}}$ position angle for central galaxies $\phi_{\mathrm{h}}$ 35.4 Normal dist. $\phi_{\mathrm{s*}}$ position angle for satellite galaxies $\phi_{\mathrm{sh}}$ 35.4 Normal dist. $\gamma_{\mathrm{ext,1}}$ external shear component 0 Eq. (22) Truncated normal dist. (Eq. (21)) $\gamma_{\mathrm{ext,2}}$ external shear component 0 Eq. (22) Truncated normal dist. (Eq. (21))
@@ 3.1 Velocity dispersion function
We calculate velocity dispersions of galaxies in our compound lens model using the GalKin package in the Lenstronomy python library (2018PDU....22..189B; 2021JOSS....6.3283B). In order to derive the velocity dispersion function of galaxies in our model, we first specify the field of view to populate halos and galaxies and then create a mock catalog of halos and galaxies in that region for each redshift bin as described in Sec. 2. We include not only main halos and central galaxies but also subhalos and satellite galaxies. In the latter case, we adopt the masses at their accretion time for computing velocity dispersions because we use those masses for calculating lensing properties of subhalos (see the discussion in Sec. 2.3 for more details). We then use the GalKin package to compute the velocity dispersion for each galaxy in the mock catalog, assuming the NFW and Hernquist profiles as density profiles of dark matter and stellar mass components, respectively. Here, the effect of the ellipticity is ignored because, as explained later, velocity dispersions are obtained from observations after circularization.
We compare the velocity dispersion function predicted by our model with that from OM10+, which is based on 2010MNRAS.404.2087B. In 2010MNRAS.404.2087B, the velocity dispersion function derived from the SDSS Data Release 6 is fitted to the Schechter function (2003ApJ...594..225S) as
@@
with $\varphi_{*}=2.099\times 10^{-2}(h/0.7)^{3}\mathrm{Mpc}^{-3},\sigma_{*}=113.78\mathrm{~km}\mathrm{~s}^{-1}$, $\alpha=0.94$, and $\beta=1.85$. OM10+ also takes account of the redshift evolution of the velocity dispersion function, mainly reflecting an evolution in the mass-to-light ratio of galaxies (2005ApJ...622L...5T; 2007ApJ...655...30V; 2010A&A...524A...6S; 2010ApJ...724..714H; 2014ApJ...793L..31V; 2015ApJ...806..122Z), as
@@
Here $\varphi_{\text{hyd }}(\sigma,z)$ is the velocity dispersion function presented in 2015MNRAS.454.2770T, which is derived from the Illustris simulation (2014MNRAS.444.1518V; 2014Natur.509..177V; 2014MNRAS.445..175G). We caution that Eqs. (23) and (24) are used only for the validation test conducted in this Section, and not used for our main calculations to generate mock catalogs based on the halo model approach.
It should be noted that this velocity dispersion function is for $\sigma_{e/8}$, i.e., a luminosity-weighted velocity dispersion averaged within the effective radius $r_{e}/8$, rather than within $r_{e}$ [^1] Conventionally, this luminosity-weighted velocity dispersion within $r=r_{e}/8$ has often been adopted, historically based on the aperture size of previous observations. In this study, since we are considering strong gravitational lensing effects and the Einstein radii are typically comparable to $r=r_{e}$ (e.g., 2014MNRAS.439.2494O), the luminosity-weighted velocity dispersion averaged within $r=r_{e}$ is more relevant than $r=r_{e}/8$.. Hence, we convert the velocity dispersion function of $\sigma_{e/8}$ to that of $\sigma_{e}$ using the commonly-used relation for elliptical galaxies, which is derived from 40 early-type galaxies of the SAURON integral field spectroscopy data in 2006MNRAS.366.1126C,
@@
Fig. 2 shows the comparison between the velocity dispersion function derived from our model and that of OM10+ calculated from Eq. (24) for several redshifts. For numerical reasons, we set a lower limit on the mass of halos to be calculated, $M_{\mathrm{hh}}\gtrsim 10^{11}~M_{\odot}$ (or $M_{\mathrm{sh}}\gtrsim 10^{10}~M_{\odot}$ for subhalos), which may lead to the underestimation of the number of galaxies with low velocity dispersions in our model. We find that both the velocity dispersion functions agree well for low redshifts ($z\lesssim 0.5$), which are main redshifts of lensing objects for lensed QSOs and SNe. In contrast, we find some mismatch at high redshifts ($z\gtrsim 0.5$), indicating that the redshift evolution in our model deviates from that in the Illustris simulation. To be more quantitative, we conduct the Kolmogorov-Smirnov test focusing on the velocity dispersion function at high velocity dispersions $v\geq 200\mathrm{~km}\mathrm{~s}^{-1}$ to find that p-values are larger than 0.002 for $z\leq 0.25$ and are smaller than $2\times 10^{-5}$ for $z\geq 0.5$, indicating that the mismatch of two velocity dispersion functions is indeed significant at high redshifts. Although the Illustris simulations were calibrated to match a set of global galaxy properties, they predicted galaxy sizes about twice as large as those observed (2014MNRAS.445..175G). In the updated IllustrisTNG simulations incorporating AGN feedback, the predicted galaxy sizes are significantly reduced (2018MNRAS.473.4077P; 2018MNRAS.474.3976G). Since there is a negative correlation between the galaxy size and the velocity dispersion for given stellar and halo masses, it is possible that the original Illustris simulations predict significantly smaller velocity dispersion at such high redshifts. As described in Sec. 2, our model adopts galaxy sizes directly measured from observations. Therefore, we expect our model to predict the redshift evolution of the velocity dispersion function more accurately and reliably than OM10+.
@@
Figure 2: Comparison between the velocity dispersion function derived with our model described in Sec. 2 (red) and that presented in 2018MNRAS.480.3842O (grey). The vertical axis shows the number of galaxies in each redshift bin with the width of $\Delta z=0.01$ for each velocity dispersion bin. Here, we set the area written on each panel to derive the number. Top left: $z=0.05$, Top center: $z=0.1$, Top right: $z=0.25$, Bottom left: $z=0.5$, Bottom center: $z=1.0$, Bottom right: $z=2.0$.
@@ 3.2 Effective radii and velocity dispersions at intermediate redshift
As mentioned above, our model of the effective radius of galaxies is based on measurements with JWST for galaxies at redshift $0.5<z<2.5$. For the intermediate redshift range of $0.2\lesssim z\lesssim 0.6$, there is a galaxy catalog from the SHELS survey of the F2 field of the Deep Lens Survey (2006ApJ...643..128W; 2014ApJS..213...35G; 2022ApJ...929...61D). As an additional validity check, we compare effective radii and velocity dispersions for early-type galaxies at $0.2\lesssim z\lesssim 0.6$ in that galaxy catalog with those predicted by our model. The four panels on the left side of Fig. 3 show the effective radii of galaxies as a function of stellar mass, focusing on massive galaxies with masses of $M_{*}\gtrsim 10^{10.7}~M_{\odot}$ at $0.2<z<0.3$, $0.3<z<0.4$, $0.4<z<0.5$, and $0.5<z<0.6$. The right panel displays the corresponding velocity dispersions. For both effective radii and velocity dispersions, the SHELS F2 catalog galaxies, especially more massive galaxies than $10^{11}~M_{\odot}$, appear to have somewhat larger value than our mock galaxies, which is, however in agreement at the $2\sigma$ level.
@@
Figure 3: Comparison between our model with observation data of SHELS F2 at intermediate redshifts, $z\lesssim 0.6$. Only the galaxy samples with stellar masses of $M_{*}\gtrsim 10^{10.7}~M_{\odot}$ are plotted. The four panels on the left side show effective radii as a function of stellar mass at $0.2<z<0.3$, $0.3<z<0.4$, $0.4<z<0.5$, and $0.5<z<0.6$, while the four panels on the right represent those of velocity dispersions. The thick solid lines show the center lines for the galaxy samples. The dark red and light red regions show the $68\%$ and $95\%$ areas for our galaxy samples, while the thin grey solid and dotted lines show the $68\%$ and $95\%$ areas for SHELS F2 galaxy samples. We set the area of $4.2~\mathrm{deg}^{2}$ to create a mock galaxy sample, matching the area covered by the SHELS F2 catalog. Note that we do not plot the error bars for the effective radii and velocity dispersions for the SHELS F2 data.
@@ 3.3 Intrinsic scatter of velocity dispersions
Next, we discuss the intrinsic scatter of velocity dispersions, $\sigma_{e}$, for a given stellar mass of galaxies, which is also important in order to create a realistic mock catalog. For instance, ignoring the intrinsic scatter typically is suggested to lead to an underestimation of the slope of the $M_{*}$-$\sigma_{e}$ relation (see, e.g., 2010ApJ...724..511A). Using the mock catalog obtained for the comparison of the velocity dispersion function, we can estimate the scatter of $\sigma_{e}$ at given $M_{*}$ and $z$. In the case of our model, the intrinsic scatter originates from three sources, (1) the scatter of the stellar-mass-halo-mass relation (2022ApJ...928...28G), (2) the scatter of the halo mass-concentration parameter relation (2007MNRAS.378...55M), and (3) the scatter of the galaxy size-mass relation (2024ApJ...960...53V). We note that 2024ApJ...960...53V claims that measurement uncertainties in observations are smaller than the intrinsic scatter of $M_{*}-r_{e}$ relation. Therefore, we assume that the scatter obtained from our model is also dominated by the intrinsic scatter.
@@
2020MNRAS.498.1101C suggest that the intrinsic scatter of the $M_{*}$-$\sigma_{e}$ relation is about $0.075_{-0.003}^{+0.003}~\mathrm{dex}$ at given $M_{*}$ for early-type galaxies with masses $M_{*}\gtrsim 3.0\times 10^{10}~M_{\odot}$ in the SDSS sample. They also investigate the redshift evolution of the intrinsic scatter using the sample of the massive early-type galaxies at higher redshifts, $z\lesssim 2$, composed of 26 galaxies drawn from the LRIS sample presented in 2014ApJ...783..117B, 56 galaxies from 2013ApJ...771...85V, four galaxies from 2015A&A...573A.110G, and 24 galaxies from 2017ApJ...834...18B. As a result, it is found that the redshift evolution of the scatter is weak in terms of the Bayesian factor comparison.
@@
We compute the intrinsic scatter in a stellar mass range of $10.5<\log_{10}(M_{*}/M_{\odot})<11.7$ at $z<2$. First, we divide our mock galaxies within this mass range into ($z,M_{*}$)-bins. We then calculate the scatter of velocity dispersion for the galaxy samples in each bin. The histograms of the scatter are plotted in Fig. 4, divided into three redshift ranges: $0.05<z<0.2$, $0.2<z<0.6$, and $0.6<z<2.0$. For the galaxy samples, we set $(\Delta z,\Delta\log_{10}(M*/M_{\odot}))=(0.01,0.02)$ for $0.05<z<0.2$, $(0.02,0.02)$ for $0.2<z<0.6$, and $(0.1,0.02)$ for $0.6<z<2.0$ for numerical reasons. It should be noted that the histograms in Fig. 4 include only the scatter from bins containing a sufficient number of galaxy samples, $N_{\mathrm{gal}}^{(z,M_{*})-\mathrm{bin}}>\mathcal{O}(10^{2})$. In the left panel of Fig. 4, we also plot the result of 2020MNRAS.498.1101C for SDSS samples containing galaxies at $0.05<z<0.2$ for comparison. We find that our model predicts the intrinsic scatter of about $0.072^{+0.002}_{-0.002}$ dex at $0.05<z<0.2$, which agrees with the result of 2020MNRAS.498.1101C at $\sim 1\sigma$ level. We also find that our model does not exhibit a strong redshift dependence on the intrinsic scatter. Since 2020MNRAS.498.1101C suggest the weak redshift evolution of the scatter, both results appear to be consistent.
@@
Figure 4: The scatter of $\sigma_{e}$ for mock galaxies at given $z$ and $M_{*}$ calculated with our model. We compute the scatter only for mock galaxies with $10.5<\log_{10}(M_{*}/M_{\odot})<11.7$ at $z<2$. We set $2000~\mathrm{deg}^{2}$, $100~\mathrm{deg}^{2}$, and $20~\mathrm{deg}^{2}$ to create mock galaxy samples at $0.05<z<0.2$ (left), $0.2<z<0.6$ (middle), and $0.6<z<2.0$ (right), respectively. The calculation method of the scatter is as follows. We first divide our mock galaxies into ($z,M_{*}$)-bins. We then calculate the scatter of velocity dispersion for the sample of galaxies in each bin. We finally plot these histograms only using those bins that hold more than $\mathcal{O}(10^{2})$ galaxy samples to ensure that the variance is properly computed. Here we set $(\Delta z,\Delta\log_{10}(M*/M_{\odot}))=(0.01,0.02)$ for the galaxy sample at $0.05<z<0.2$, $(0.02,0.02)$ at $0.2<z<0.6$, and $(0.1,0.02)$ at $0.6<z<2.0$ for a numerical reason. The grey dotted line appearing only in the left panel represents a best-fit value obtained for the SDSS galaxy sample at $0.05<z<0.2$ in 2020MNRAS.498.1101C, $\sigma_{\ln\sigma_{e}}=0.075^{+0.003}_{-0.003}~\mathrm{dex}$. The orange-colored regions represent the $\pm 1\sigma$, $\pm 2\sigma$, and $\pm 3\sigma$ regions, in order from darkest to lightest.